# Diffusion Model with Perceptual Loss

### v6

@@ Abstract
Diffusion models without guidance tend to generate unrealistic samples, yet the cause of this problem is not fully studied. Our analysis suggests that the loss objective plays an important role in shaping the learned distribution and the common mean squared error loss is not optimal. We hypothesize that a better loss objective can be designed with inductive biases and propose a novel self-perceptual loss that utilizes the diffusion model itself as the perceptual loss. Our work demonstrates that perceptual loss can be used in diffusion training to improve sample quality effectively. Models trained using our objective can generate realistic samples without guidance. We hope our work paves the way for more future explorations of the diffusion loss objective.
@@ 1 Introduction
Conceptually, diffusion models [15, 47, 50] work by transforming pure noise to data samples through repeated denoising. Formally, each denoising step can be viewed from the lens of score matching [50] such that the model learns to predict the drift (score) of a stochastic differential equation (SDE), or equivalently the gradient (flow) of a ordinary differential equation (ODE), that transports samples from one distribution (noise) to another distribution (image, video, etc.) [29]. In this paper, our experiments focus on image generation.
@@
Although diffusion models are supposed to transport samples from the noise to the image distribution by theory, images generated by diffusion models without guidance are often of poor quality as shown in LABEL:fig:banner, despite the improvements in model architecture [4, 21, 34, 36, 41, 38, 39, 8, 28], formulation [27, 29, 20], and sampling strategy [48, 30, 31, 20].
Classifier-free guidance (CFG) [14] is applied almost ubiquitously in state-of-the-art diffusion models across modalities to improve sample quality, e.g., text-to-image [36, 39, 41, 4, 38, 55, 35], text-to-video [11, 2, 56, 16, 46, 1], text-to-3d [45, 52, 37], image-to-video [1, 53], video-to-video [3, 7], etc. However, prior works have shown that CFG reduces mode coverage [14, 22] and can generate out-of-distribution samples (over-exposed or over-saturated images) [25, 41, 24], etc.
In this paper, we seek to study the fundamental question of why diffusion models without guidance fail to generate desirable samples. Our analysis suggests that the loss objective has an important role in shaping the learned distribution, and we hypothesize that the common squared distance loss objective is not optimal. Based on these findings, we experiment using a perceptual distance loss objective. Specifically, we propose a novel self-perceptual objective that uses the diffusion model itself as the perceptual loss. Our method generates visibly more realistic samples without guidance. Our main contributions are as follows:
  • We analyze the effect of loss objective in shaping the learned probability distribution of diffusion models, which leads to our discovery that the squared distance objective is causing poor generation quality.
  • To the best of our knowledge, we are the first to apply perceptual loss to diffusion training. We demonstrate its effectiveness in improving sample quality while having good mode coverage.
We want to emphasize that our contributions are beyond the proposal of the perceptual objective. Rather, our work studies the underlying cause of why diffusion models generate poor samples without guidance. Our analysis and experiment show a better loss objective can address the issue and pave the way for more future explorations in this area.
@@ 2 Related Work
Guidance methods alter the model prediction and guide the sample toward desired regions during the generation process. Classifier Guidance [6] adds classifier gradients to the predicted score to guide the sample generation to maximize the classification. It can turn an unconditional diffusion model conditional. However, it is not evident why applying classifier guidance on an already conditional diffusion model can further improve sample quality. Classifier-Free Guidance [14] uses Bayes’ rule and finds that the diffusion model itself can be inverted as an implicit classifier. Specifically, the model is queried both conditionally and unconditionally at every inference step and the difference is amplified toward the conditional direction. However, both methods only work for conditional generation and entangle sample quality with conditional alignment [22]. More recently, Discriminator Guidance [23] proposes to train a discriminator network to classify real and generated samples and uses it as guidance during diffusion generation. Self-Attention Guidance [17] finds that the self-attention map of the diffusion model can be exploited to enhance quality. Autoguidance [22] finds a smaller or less-trained model can be used as negative guidance to improve quality. Although these methods are effective, why diffusion models generate poor samples without guidance is not addressed.
The loss objective of diffusion models has been studied by prior works. Loss weighting is found to influence perceptual quality and likelihood evaluation [5, 49, 12] but still cannot produce good samples without guidance. Multi-scale loss [18] is proposed to improve high-resolution generation. Smoothness penalty [10] is proposed to enforce smoother latent traversal. $l1$ distance is explored for colorization and in-painting but reduces diversity [40]. The squared distance objective is almost ubiquitously adopted.
In recent years, the size of diffusion models has increased to multi-billion parameters [8, 28]. The architecture has improved from convolution [39, 36, 41] to transformers [8, 28, 4]. More accurate solvers [30, 31] and better formulations [29, 27, 20] have been proposed. However, diffusion models still generate poor samples without guidance. This leads us to study the underlying cause.
@@ 3 Analysis
In this section, we analyze why diffusion models without guidance generate poor samples. We hypothesize that the loss objective is important in shaping the learned distribution and the squared distance is not optimal. This inspires the exploration of perceptual loss to address the issue.
@@ 3.1 The Effect of the Loss Objective
A reasonable assumption is to model the distribution with high probabilities over the observed samples. A special solution, the maximum likelihood estimation (MLE) distribution, only assigns probability over the observed samples and zero probability everywhere else. Its density function can be expressed as a sum of Dirac delta functions as illustrated in Fig. 1(b). Such distribution overfits and only reproduces the observed samples at test time.
For diffusion training, the target probability flow is unique and pre-determined by the forward diffusion process [50]. It transports between the noise distribution and the MLE data distribution [50, 49]. This means that with sufficient model capacity, the model will learn the MLE data distribution and overfit to only generate the observed samples. In practice, models are parameterized by neural networks with finite capacity, which smooths the prediction and provides desired generalization, resulting in distributions like Figs. 1(c) and 1(d).
The shape of the learned distribution can be influenced by the model capacity, architecture, and the loss objective. In particular, we hypothesize that the loss objective can be better designed with inductive biases to drive the shape of the learned distribution toward the actual distribution.
@@ 3.2 The Problem of the Squared Distance
The squared distance was originally proposed from theoretical derivations. From the diffusion theory perspective, the reverse generative process has the same functional form as the forward Gaussian diffusion process [47]. KL divergence is chosen to measure the Gaussian divergence between the model prediction and the posterior ground truth [15]. The negative log-likelihood for Gaussians can be further simplified to the squared distance. From the score matching perspective [19] and the most recent flow matching perspective [27], the squared distance seems to be chosen out of convenience.
We argue that KL divergence is not the only divergence measurement and simply extending the theory to high-dimensional data (images) assumes the independence of each dimension (pixels), which is generally not true.
Figure 2 illustrates that the per-pixel squared distance (MSE) objective is a poor metric for measuring the distance between images. Midpoint samples derived by minimizing the distance to known samples are completely out of distribution. This leads us to hypothesize that the squared distance may be the cause for diffusion models generating out-of-distribution samples.
Figure 2: Squared distance assumes the independence of each data dimension and is generally not a good distance measurement for data distributions such as images.
@@ 3.3 The Effect of the Guidance
This may explain the effectiveness of guidance in improving diffusion generation quality. In the case of classifier guidance [6], the classifier network captures the image semantics and guides the generation toward perceptually realistic samples as a side effect of maximizing the classification. Classifier-free guidance [14], derived using Bayes’ rule and using the diffusion model itself as an implicit classifier, may be explained for the same reason. Additionally, Generative Adversarial Networks (GANs) [9] do not exhibit the pathology because the discriminator, being a deep neural network, better captures the data semantics.
However, guidance has limitations, e.g., reduction in mode coverage [14, 22], out-of-distribution generation manifested as over-exposure or over-saturation in images [25, 41, 24], and additional computation at inference.
The above analysis inspires us to explore substituting the squared distance with a better-designed distance function, such as the perceptual distance, to directly address the issue.
@@ 4 Method
In this section, we describe our method to incorporate perceptual distance loss into diffusion training. In Sec. 4.1 we introduce the diffusion background and our model formulation. In Sec. 4.2, we propose a novel self-perceptual objective and show that the diffusion model itself can be used as a perceptual network to provide meaningful perceptual loss.
@@ 4.1 Background
The original squared distance objective can be written as the following MSE loss:
@@ 4.2 Self-Perceptual Objective
Prior research has shown that the distance computed on the feature maps of a pre-trained deep neural network is a good perceptual metric [54].
We find that the diffusion model trained with the vanilla MSE loss is effectively a perceptual network. The score matching objective is strongly related to the denoising autoencoder objective, which can train the model to learn data semantics [51]. This should not come as a surprise as the classifier-free guidance [14] has already successfully exploited the diffusion model itself to provide meaning perceptual guidance at inference.
Our approach is to employ the MSE pre-trained diffusion model as our perceptual network and compute distance on its feature maps. The model is suitable for our task because it is trained on the target dataset, on the latent space, and on all noise levels $x_{t}$.
First, we copy and freeze the diffusion model trained with the MSE loss, and we modify the architecture to return the hidden feature at layer $l$. We denote this frozen perceptual network as $p^{l}_{*}$.
@@
Notice that the $\hat{x}_{t}^{\prime}$ derived here is equivalent to a single Euler solver step from timestep $t$ to $t^{\prime}$.
@@
$\mathcal{L}_{sp}=\|p^{l}_{*}(\hat{x}_{t^{\prime}},t^{\prime},c)-p^{l}_{*}(x_{t^{\prime}},t^{\prime},c)\|_{2}^{2}.$ (11)
The pseudo-code is provided in the supplementary materials.
@@ 5 Evaluation
We first finetune Stable Diffusion v2.1 [39] using our formulation and MSE loss $\mathcal{L}_{mse}$ on a subset of LAION dataset [44] for 60k iterations. This ensures that the training data is consistent for fair comparisons. We use 10% conditional dropout to support CFG for evaluation comparison. Then we copy and freeze the MSE model as our perceptual network, and continue training the online network with our self-perceptual objective $\mathcal{L}_{sp}$ for 50k iterations. We use learning rate 3e-5, batch size 896, and EMA decay 0.9995.
@@ 5.1 Qualitative
Figure 3 shows the conditional generation results. Our self-perceptual objective has significant quality improvement over the MSE objective. This validates our analysis in Secs. 3.1 and 3.2 that the squared distance is the cause for poor generation quality and a better loss objective such as the perceptual distance can indeed improve quality.
Notice that the results of the MSE and the self-perceptual objective share very similar content and layout when generated from the same initial noise. As stated in Sec. 3.1, this is because the underlying target MLE probability flow is exactly the same, and changing the loss objective only influences the model’s learned generalization, so the same noise will map to data in the similar region. Furthermore, the learned distribution should generally reflect and capture all the modes in the MLE distribution under a reasonable choice of the distance function, so our approach should maintain good mode coverage. On the other hand, CFG does not have this property and is prone to mode dropping.
Compared to CFG, the self-perceptual objective only affects sample quality but not text alignment. This is especially evident in Fig. 3(j), where CFG enhances the text condition. Since the LAION dataset contains dirty image captions, this gives CFG an unfair advantage.
@@ 5.2 Quantitative
Table 1 shows the quantitative evaluation for conditional generation. We follow the convention to calculate Fréchet Inception Distance (FID) [13, 33] and Inception Score (IS) [43]. We select the first 10k samples from the MSCOCO 2014 validation dataset [26] and use our models to generate images of the corresponding captions. Our self-perceptual objective has improved FID/IS over the vanilla MSE objective. This aligns with the improvement observed in our qualitative comparison. However, CFG still achieves better scores than our self-perceptual objective. We suspect this may be due to LAION data containing dirty captions which gives CFG an advantage, and there are additional architectural optimizations we can improve as discussed later in the ablation studies. Despite this, the result is sufficient to support the main claim that the loss objective is important in shaping the learned distribution and our results demonstrate that the perceptual loss can improve sample quality.
@@
$\mathcal{L}_{sp}$ 25 25 25.89 27.76
50 50 24.42 28.07
Table 1: Conditional generation. Quantitative evaluation on MSCOCO 10K validation dataset. Our self-perceptual (SP) objective improves FID and IS metrics over vanilla MSE objective but has not surpassed classifier-free guidance [14] with rescale [25]. Since classifier-free guidance with 25 steps incurs 50 NFEs (number of function evaluations), we show both 25-step and 50-step metrics.
Loss IS $\uparrow$
$\mathcal{L}_{mse}$ 11.18
$\mathcal{L}_{sp}$ 12.04
Table 2: Unconditional generation metrics. The self-perceptual objective improves IS over the vanilla MSE objective.
Table 2 shows that our approach also improves IS for unconditional generation. Since we cannot control the generated content through captions, the FID comparison against MSCOCO is not calculated.

### v7

@@ Abstract
Diffusion models without guidance generate very unrealistic samples. Guidance is used ubiquitously, and previous research has attributed its effect to low-temperature sampling that improves quality by trading off diversity. However, this perspective is incomplete. Our research shows that the choice of the loss objective is the underlying reason raw diffusion models fail to generate desirable samples. In this paper, (1) our analysis shows that the loss objective plays an important role in shaping the learned distribution and the MSE loss derived from theories holds assumptions that misalign with data in practice; (2) we explain the effectiveness of guidance methods from a new perspective of perceptual supervision; (3) we validate our hypothesis by training a diffusion model with a novel self-perceptual loss objective and obtaining much more realistic samples without the need for guidance. We hope our work paves the way for future explorations of the diffusion loss objective.
@@ 1 Introduction
Conceptually, diffusion models [16, 48, 51] work by transforming noise to data samples through repeated denoising. Formally, each denoising step can be viewed from the lens of score matching [51] such that the model learns to predict the drift (score) of a stochastic differential equation (SDE), or equivalently the gradient (flow) of an ordinary differential equation (ODE), that transports samples from one distribution (noise) to another distribution (image, video, etc.) [30].
@@
Although diffusion models are supposed to transport samples from the noise to the data distribution by theory, samples generated by diffusion models without guidance are often of poor quality as shown in LABEL:fig:banner, despite the improvements in model architecture [5, 23, 35, 37, 42, 39, 40, 9, 29], formulation [28, 30, 22], and sampling strategy [49, 31, 32, 22].
Classifier-free guidance (CFG) [15] is applied almost ubiquitously in state-of-the-art diffusion models across modalities to improve sample quality, e.g., text-to-image [37, 40, 42, 5, 39, 56, 36], text-to-video [12, 2, 57, 17, 47, 1], text-to-3d [46, 53, 38], image-to-video [1, 54], video-to-video [3, 8], etc. Previous research has attributed its effect to low-temperature sampling [15], as if the quality improvement is a result of trading off diversity. However, our research provides a different perspective.
In this paper, we seek to uncover the fundamental cause of why diffusion models without guidance fail to generate desirable samples. Our analysis suggests that the loss objective plays an important role in shaping the learned distribution, and the common MSE loss objective derived from theories holds assumptions that misalign with data in practice. Based on these findings, we experiment using a perceptual distance loss objective. Specifically, we propose a novel self-perceptual objective that uses the diffusion model itself as the perceptual loss. Our method generates much more realistic samples without guidance. Our main contributions are as follows:
  • Our analysis uncovers the important effect of the loss objective in shaping the learned probability distribution of diffusion models, and shows that the MSE loss holds assumptions that misalign with data in practice. More importantly, we show that the loss objective is open for exploration without a single theoretically correct solution.
  • We provide a different perspective on guidance methods through the lens of perceptual supervision instead of low-temperature sampling.
  • To the best of our knowledge, we are the first to apply perceptual loss to diffusion training. We propose a novel self-perceptual loss that uses the diffusion model itself as the perceptual network. We demonstrate its effectiveness in improving sample quality.
Our work studies the underlying cause of why diffusion models generate poor samples without guidance. We hope our work paves the way for more future explorations in the diffusion loss objective.
@@ 2 Related Work
Guidance methods alter the model prediction and guide the sample toward desired regions during the generation process. Classifier Guidance [7] adds classifier gradients to the predicted score to guide the sample generation to maximize the classification. It can turn an unconditional diffusion model conditional. However, it is not evident why applying classifier guidance on an already conditional diffusion model can significantly improve sample quality. Previous research has attributed it to low-temperature sampling [15, 24]. Classifier-Free Guidance (CFG) [15] uses Bayes’ rule and finds that the diffusion model itself can be inverted as an implicit classifier. Specifically, the model is queried both conditionally and unconditionally at every inference step and the difference is amplified toward the conditional direction. Both methods only work for conditional generation and entangle sample quality with conditional alignment [24]. Self-Supervised Guidance [20] uses self-supervised networks to generate synthetic clustering labels for unconditional data. This allows unconditional data to use CFG for improving quality. Guidance-Free Training [4] shows that CFG can be applied at training. This bakes the guided flow into the model and saves computation during inference. More recently, Discriminator Guidance [25] proposes to train a discriminator network to classify real and generated samples and use it as guidance during diffusion generation. Self-Attention Guidance [18] finds that the self-attention map of the diffusion model can be exploited to enhance quality. Autoguidance [24] finds a smaller or less-trained model can be used as negative guidance to improve quality. Although these methods are effective in improving quality, they also present issues such as increased complexity and reduced diversity [15, 24], etc. On the other hand, our research aims to explore the underlying issue: why diffusion models without guidance fail in the first place.
The loss objective of diffusion models has been studied by prior works. Loss weighting is found to influence perceptual quality and likelihood evaluation [6, 50, 13] but still cannot produce good samples without guidance. Multi-scale loss [19] is proposed to improve high-resolution generation. Smoothness penalty [11] is proposed to enforce smoother latent traversal. $l1$ distance is explored for colorization and in-painting tasks [41]. The squared distance objective is almost ubiquitously adopted. Perceptual loss has been explored in consistency models [52], but we are the first to explore perceptual loss in diffusion models.
In recent years, the size of diffusion models has increased to multi-billion parameters [9, 29]. The architecture has improved from convolution [40, 37, 42] to transformers [9, 29, 5]. More accurate solvers [31, 32] and better formulations [30, 28, 22] have been proposed. However, diffusion models still generate poor samples without guidance.
@@ 3 Analysis
In this section, we analyze why diffusion models without guidance generate poor samples. We show that the loss objective is important in shaping the learned distribution and the MSE loss is not optimal.
@@ 3.1 The Effect of the Loss Objective
A special solution, the maximum likelihood estimation (MLE) distribution, only assigns probability over the observed samples and zero probability everywhere else. Its density function can be expressed as a sum of Dirac delta functions as illustrated in Fig. 1(b). Notice that the MLE distribution is always spiky regardless of the dataset size as long as the samples are finite because having more samples will just result in closer spikes. The MLE distribution overfits and only reproduces the observed samples at test time.
For diffusion training, the target probability flow is unique and pre-determined by the forward diffusion process [51]. The target flow transports between the noise distribution and the MLE data distribution [51, 50]. This means that with sufficient model capacity, the model will learn the MLE data distribution and overfit to only generate the observed samples. In practice, models are parameterized by neural networks with finite capacity, which smooths the prediction and provides desired generalization, resulting in distributions like Figs. 1(c) and 1(d).
The shape of the learned distribution are influenced by the model capacity and the loss objective. In particular, the loss objective can be better designed with inductive biases to drive the shape of the learned distribution toward the actual distribution. This is often overlooked by existing research.
@@ 3.2 The Problem of the Squared Distance
The squared distance was originally proposed from theoretical derivations. From the diffusion theory perspective, the reverse generative process has the same functional form as the forward Gaussian diffusion process [48]. KL divergence was chosen to measure the Gaussian divergence between the model prediction and the posterior ground truth [16]. The negative log-likelihood for Gaussians can be further simplified to the squared distance. From the score matching perspective [21] and the most recent flow matching perspective [28], the squared distance seems to be chosen out of convenience.
First, KL divergence is not the only valid choice of divergence measurement, so the MSE loss is not the only valid choice from the theoretical viewpoint. Second, simply extending it to high-dimensional data (images) assumes the independence of each dimension (pixels), which is generally not true.
Figure 2 illustrates that the per-pixel MSE objective is a poor distance function for images. For example, given finite images of human faces, the underlying distribution is ambiguous to the model. The loss objective dictates the shape of the learned distribution. We would like the model to produce a distribution of semantic morphing of new faces, but MSE leads the model to learn a distribution of pixel-wise blending. This is why diffusion models with MSE loss produce out-of-distribution samples.
Even if the diffusion models employ a convolution or attention architecture that takes all the pixels into consideration, this only helps the model to predict more accurate pixel-independent MSE midpoints because it is trained to do so. The MSE objective is problematic in the first place.
Figure 2: The midpoint sample is derived by minimizing the distance to known samples by the given distance function. MSE midpoint is out-of-distribution.
@@ 3.3 The Effect of the Guidance
This explains the effectiveness of guidance in improving diffusion generation quality. In the case of classifier guidance [7], the classifier network captures the image semantics and guides the generation toward perceptually realistic samples as a side effect of maximizing the classification. For example, when the classifier is asked to classify human faces, it will assign high scores to samples with semantically correct faces with only two eyes and will penalize the pixel-blending samples that may have four eyes. Classifier-free guidance [15], derived using Bayes’ rule and using the diffusion model itself as an implicit classifier, can explained for the same reason. Additionally, Generative Adversarial Networks (GANs) [10] do not exhibit the pathology because the discriminator, being a deep neural network, learns to better capture the data semantics. Discriminator guidance can be explained in the same way.
We do believe low-temperature sampling is indeed a factor, as maximizing the classifier score eliminates the low-likelihood samples. However, in our perspective, we believe the MSE loss objective is the main reason diffusion models fail to generate desirable distributions without guidance, and guidance methods provide perceptual supervision in the sampling process.
@@ 4 Method
To validate our analysis, we experiment with directly incorporating perceptual loss into diffusion training. In Sec. 4.1 we introduce the diffusion background and our model formulation. In Sec. 4.2, we propose a novel self-perceptual objective and show that the diffusion model itself can be used as a perceptual network to provide meaningful perceptual loss.
@@ 4.1 Background
The original MSE objective is defined as:
@@ 4.2 Self-Perceptual Objective
We propose a self-perceptual objective that utilizes the MSE diffusion model itself as the perceptual network. This is not surprising as the classifier-free guidance [15] also exploits the MSE diffusion model itself to provide meaning perceptual guidance at inference.
The intuition is that even though the model’s MSE flow prediction is not ideal, the model is still trained with good semantic understanding in order to predict accurate MSE flow. Therefore, our approach employs the MSE pre-trained diffusion model as our perceptual network and computes distance on its feature maps following prior perceptual loss research [55].
Specifically, we copy and freeze the diffusion model trained with the MSE loss, and we modify the architecture to return the hidden feature at layer $l$. We denote this frozen perceptual network as $p^{l}_{*}$, and the online trainable network as $f_{\theta}$.
@@
Notice that the $\hat{x}_{t}^{\prime}$ derived here is equivalent to a single DDIM solver step from timestep $t$ to $t^{\prime}$.
@@
$\mathcal{L}_{sp}=\|p^{l}_{*}(\hat{x}_{t^{\prime}},t^{\prime},c)-p^{l}_{*}(x_{t^{\prime}},t^{\prime},c)\|_{2}^{2},$ (11)
The code is provided in the supplementary materials.
We highlight that this design intentionally avoids introducing any external components, allowing us to isolate the study on loss objective. It is also practical, as we can easily finetune existing MSE diffusion models to SP using itself.
@@ 5 Evaluation
We first finetune Stable Diffusion v2.1 [40] using our formulation and MSE loss $\mathcal{L}_{mse}$ on a LAION aesthetic 6+ dataset [45] for 60k iterations. This ensures that the training data is consistent for fair comparisons. We use 10% conditional dropout to support CFG for evaluation comparison. Then we copy and freeze the MSE model as our perceptual network, and continue training the online network with our self-perceptual objective $\mathcal{L}_{sp}$ for 50k iterations. We use learning rate 3e-5, batch size 896, and EMA decay 0.9995. We verify both networks are trained till convergence for a fair comparison.
@@ 5.1 Qualitative
Figure 3 shows the conditional generation results. Our self-perceptual objective has significant quality improvement over the MSE objective. This validates our analysis in Secs. 3.1 and 3.2 that the MSE loss is the cause for poor generation quality and a better loss objective such as the perceptual distance can indeed improve quality.
Notice that the results of the MSE and the self-perceptual objective share very similar content and layout when generated from the same initial noise. As stated in Sec. 3.1, this is because the underlying target MLE probability flow is exactly the same, and changing the loss objective only influences the model’s learned generalization, so the same noise will map to data in the similar region. On the other hand, CFG changes the flow drastically.
Compared to CFG, the self-perceptual objective only affects sample quality but not text alignment. This is especially evident in Fig. 3(j), where CFG enhances the text condition, while our SP model and the original MSE model are more aligned with the training dataset distribution, where LAION dataset has less-aligned dirty captions.
@@
Figure 3: Text-to-image generation on DrawBench prompts [42]. Our self-perceptual objective improves sample quality over the MSE objective while largely maintaining the image content and layout. Classifier-free guidance has the additional effect of enhancing text alignment by sacrificing sample diversity. Images are generated with DDIM 50 NFEs. More analysis in Sec. 5.1.
@@ 5.2 Quantitative
Table 1 shows the quantitative evaluation for conditional generation. We follow the convention to calculate Fréchet Inception Distance (FID) [14, 34] and Inception Score (IS) [44]. We select the first 10k samples from the COCO 2014 validation dataset [27] and use our models to generate images of the corresponding captions.
Our self-perceptual objective has significantly improved FID/IS over the vanilla MSE objective. This aligns with the improvement observed in our qualitative comparison and validates our analysis. Additionally, we show comparisons with CFG Our SP objective is still weaker compared to CFG. Specifically, we are close to CFG in FID, but CFG+Rescale [26] is still better in both metrics. However, we emphasize that our main comparison target is the MSE baseline for guidance-less generation. Our research focus is to demonstrate the effect of the loss objective instead of to claim the new state-of-the-art. We discuss limitations and future improvements in Sec. 6.7.
@@
Table 1: Conditional generation. Quantitative evaluation on MSCOCO 10K validation dataset. Our self-perceptual (SP) objective improves FID and IS metrics over MSE objective but has not surpassed classifier-free guidance [15] with rescale [26]. Since classifier-free guidance with 25 steps incurs 50 NFEs (number of function evaluations), we show both 25-step and 50-step metrics.
Loss FID $\downarrow$ IS $\uparrow$
$\mathcal{L}_{mse}$ 62.32 11.18
$\mathcal{L}_{sp}$ 59.12 12.04
Table 2: Unconditional generation metrics. The self-perceptual objective improves FID and IS over the MSE objective.
Table 2 shows that our approach also improves FID/IS for unconditional generation.